# Learning solution operator of dynamical systems with diffusion maps kernel ridge regression

### v1

@@ Abstract
Many scientific and engineering systems exhibit complex nonlinear dynamics that are difficult to predict accurately over long time horizons. Although data-driven models have shown promise, their performance often deteriorates when the geometric structures governing long-term behavior are unknown or poorly represented. We demonstrate that a simple kernel ridge regression (KRR) framework, when combined with a dynamics-aware validation strategy, provides a strong baseline for long-term prediction of complex dynamical systems. By employing a data-driven kernel derived from diffusion maps, the proposed Diffusion Maps Kernel Ridge Regression (DM-KRR) method implicitly adapts to the intrinsic geometry of the system’s invariant set, without requiring explicit manifold reconstruction or attractor modeling, procedures that often limit predictive performance. Across a broad range of systems, including smooth manifolds, chaotic attractors, and high-dimensional spatiotemporal flows, DM-KRR consistently outperforms state-of-the-art random feature, neural-network and operator-learning methods in both accuracy and data efficiency. These findings underscore that long-term predictive skill depends not only on model expressiveness, but critically on respecting the geometric constraints encoded in the data through dynamically consistent model selection. Together, simplicity, geometry awareness, and strong empirical performance point to a promising path for reliable and efficient learning of complex dynamical systems.
@@ 1 Introduction
Broadly speaking, there are two common approaches in literature. The fully data-driven approach ignores the first-principle model and learns the entire solution operator or the vector field of the ODEs directly from observed data (see, e.g., [wang2005gaussian, williams2015data, brunton2016discovering, vlachas2018data, pathak2018model, hamzi2021learning, huang2025learning]). In contrast, the partially data-driven approach retains the first-principles model and learns only the model-error components that correct the underlying dynamics (see e.g.,[kaheman2019learning, lin2021data, harlim2021machine, levine2022framework]). In this study, we focus on the fully data-driven setting, although the proposed framework could, in principle, also be adapted for learning model-error components in partially data-driven formulations.
@@
Many methods have been proposed to identify low-dimensional representations of dynamical systems, ranging from linear techniques such as PCA (Principal Component Analysis) [jolliffe2011principal] and DMD (Dynamic Mode Decomposition) [williams2015data] and its variants, to nonlinear approaches based on autoencoders [Linot2020, Maulik2021, vlachas2018data, Regazzoni2024]. A limitation of PCA and DMD is that they sometimes fail to produce a finite, low-dimensional representation capable of capturing the full dynamics, even when the underlying invariant set itself is finite-dimensional. Moreover, algorithms in the DMD family approximate the Koopman operator and therefore require additional treatment to avoid spectral pollution. A promising algorithm to avoid spectral pollution in the Koopman spectral approximation is ResDMD [colbrook2024rigorous], which we include as a baseline in one of our test cases. As for autoencoders, it is found that the latent space only provides a global parametrization of the invariant set, whose dimension is typically larger than the intrinsic dimension of the invariant set, and the dynamics are not guaranteed to learn the exponential maps (for discrete-time model) or vector fields (for continuous-time model) while satisfying the invariant set constraint [huang2025learning]. As a result, the predicted trajectories may deviate from the invariant set, limiting the long-term prediction accuracy, as we shall see in several numerical examples with neural network (NN) baselines. Meanwhile, there are indeed autoencoder-based approaches that can uncover the invariant set when it is a smooth manifold [Floryan2022]. In such a case, the estimated dynamics are guaranteed to evolve on the manifold. However, such approach cannot be extended to the case where the invariant set is fractal. In addition, common to NN-based approaches, the computational effort to train the models is not negligible and the hyperparameter tuning process can be difficult.
@@
A promising approach with solid theoretical foundations for mitigating the curse of dimensionality is to exploit the “kernel trick,” which circumvents the explicit evaluation of high-dimensional feature spaces. For dynamical systems evolving on smooth manifolds (i.e., those whose forward invariant sets form smooth submanifolds of the phase space), the Geometric Constrained Multivariate Kernel Ridge Regression (GMKRR) method was recently introduced to approximate the governing vector field of the ODEs [huang2025learning]. GMKRR is paired with a Normal-Correction scheme that ensures time integration remains consistent with the manifold constraint. Benchmark studies have shown that GMKRR outperforms several state-of-the-art NN-based methods [Floryan2022, Regazzoni2024], achieving superior long-term predictive accuracy while reducing training costs by 3–4 orders of magnitude. Despite these promising results, GMKRR remains limited by the accuracy of tangent-bundle approximation and cannot be directly applied to systems whose invariant sets exhibit fractal structure.
@@
$x_{j}(t_{n+1})=\varphi_{j,\Delta t}({\bf x}(t_{n}))\approx f_{j}({\bf x}(t_{n}))=k_{\epsilon,N}({\bf x}_{n},X)\boldsymbol{\alpha}_{j}$ (1)
@@
$x_{j}(t_{n+1})-x_{j}(t_{n})=\varphi_{j,\Delta t}({\bf x}(t_{n}))-x_{j}(t_{n})\approx f_{j}({\bf x}(t_{n}))=k_{\epsilon,N}({\bf x}_{t},X)\boldsymbol{\alpha}_{j},$ (2)
@@
for each component $j=1,\ldots,n$. Here, $k_{\epsilon,N}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ denotes the data-dependent DM kernel with a lengthscale parameter $\epsilon$, $X=\{{\bf x}_{n}={\bf x}(t_{n})=(x_{1}(t_{n}),\ldots,x_{n}(t_{n}))\}_{n=1,\ldots,N}$ denotes the training dataset with $t_{n}=n\Delta t$, and $\boldsymbol{\alpha}_{j}$ denotes the $N$-dimensional regression coefficient for the $j$th component. While direct estimator can be used in general, we notice substantial improvement with skip-connection estimator when the underlying ODE systems are not numerically stiff. In this study, we consider a Kernel Ridge Regression solution with DM kernel, coined as Diffusion Maps Kernel Ridge Regression (DM-KRR), where the regression coefficient, $\boldsymbol{\alpha}_{j}=\boldsymbol{\alpha}_{j}(\epsilon,{\lambda_{\mathrm{reg}}})$, depends on the shape parameter, $\epsilon$, and ridge regression parameter, $\lambda_{\mathrm{reg}}$} (see the Method section for the detailed formulation). We will show that careful choice of validation metric to determine the model hyperparameters $(\epsilon,{\lambda_{\mathrm{reg}}})$ is essential for achieving optimal predictive performance. While the conclusion here resonates with the conclusion from the Kernel Flow (KF) approach [hamzi2021learning], the present approach differs significantly from KF. Particularly, KF parameterizes the flow map using a family of kernels and selects among them via validation over the same loss-square error, we employ a fixed, data-driven kernel derived the from diffusion maps and validate by comparing the prediction error of iterating the approximate model over a short time window.
@@
Beyond its simplicity and strong empirical performance, our use of kernel methods is also theoretically motivated. If the underlying dynamical system is ergodic with invariant measure $\mu$ supported on a compact set $A\subset\mathcal{M}\subset\mathbb{R}^{n}$, then the Reproducing Kernel Hilbert Space (RKHS) induced by any continuous kernel $k$ on a compact set $A$ is universal for approximating $L^{2}(\mu)$ functions, even when $A$ lacks smoothness. Thus, in principle, one may employ any such kernel, e.g., Gaussian or Matérn. Additionally, when the forward invariant set $\mathcal{M}\supset A$ is a smooth $d$-dimensional sub-Riemannian manifold and the invariant measure is uniformly distributed with a compact support that lies in the invariant set, the diffusion maps (DM) kernel is a natural choice as it approximates the heat kernel in $L^{\infty}$-sense over the dataset. In effect, the DM kernel encodes the intrinsic geometry of the invariant set, allowing the learned dynamics to adapt seamlessly to its true intrinsic dimensionality.
In the next section, we demonstrate the advantages of the DM kernel over Gaussian kernels across a range of dynamical systems whose forward invariant sets may be smooth or non-smooth and are embedded in ambient spaces ranging from dimension $3$ to $10^{5}$. In several examples, we also compare against autoencoder-based approaches and ResDMD.
2 Results
2.1 Summary of Key Findings
First, we find that, using validation with an appropriate, dynamics-aware metric, the kernel ridge regression (KRR) formulation using a scalar-valued kernel is remarkably effective across manifold, chaotic, and high-dimensional dynamical systems, outperforming state-of-the-art random feature, neural network, and operator-learning baselines. Second, within the KRR family, the diffusion maps (DM) kernel consistently delivers higher accuracy and greater sample efficiency than the conventional Gaussian RBF kernel, owing to its ability to encode the intrinsic geometry of the forward invariant set. In sum, a simple KRR framework with appropriate validation already surpasses modern alternatives, and DM further strengthens performance by adapting to the data geometry.
2.2 Evaluation Metrics
For a trajectory of length $T$, $\{{\bf x}_{i}\}_{i=1}^{T}$, suppose the rollout prediction is $\{\hat{{\bf x}}_{i}\}_{i=1}^{T}$, where $\hat{{\bf x}}_{i+1}={\bf f}(\hat{{\bf x}}_{i})$ for $i=1,2,\cdots,T-1$, with $\hat{{\bf x}}_{1}={\bf x}_{1}$ and $\bf{f}$ denoting the learned solution operator. The evaluation metrics are defined as follows,
  1. 1.
    Root Mean Squared Error (RMSE): The RMSE is evaluated over the prediction horizon as
    $\text{RMSE}=\sqrt{\frac{1}{Tn}\sum_{i=1}^{T}\left\|\hat{{\bf x}}_{i}-{\bf x}_{i}\right\|_{2}^{2}}.$ (3)
  2. 2.
    Valid Prediction Time (VPT): The VPT is defined as the maximum duration for which the normalized trajectory error remains below a certain error tolerance $\gamma$,
    $\mathrm{VPT}=\frac{1}{T_{\Lambda}}\max\left\{i|E_{i}\leq\gamma\text{ for all }i\leq T\right\},$ (4)
2.3 Dynamics on Torus
$\dot{z}_{1}$ $=c_{1}z_{2}z_{3}$ (5a)
$\dot{z}_{2}$ $=c_{2}z_{1}z_{3}$ (5b)
$\dot{z}_{3}$ $=c_{3}z_{1}z_{2},$ (5c)
where $c_{1}=1/2$, $c_{2}=-7/8$, and $c_{3}=3/8$, so that the trajectories with initial conditions on the unit sphere $S^{2}$ remain on $S^{2}$ for all time, embedded in $\mathbb{R}^{3}$. To evaluate the performance under various ambient dimension, $n$, we map the coordinates $z_{1},z_{2},z_{3}]$ on $S^{2}$ to coordinates $\bf x$} on tori embedded in $\mathbb{R}^{n}$ for various dimensions $n\in\{3,7,15\}$. However, regardless of the ambient dimension $n$, the forward invariant set of the dynamics is diffeomorphic to three quarters of $S^{2}$. See SI A.1 for more details on the coordinate transformation.
For training, $100$ initial conditions are sampled on a uniform grid on $S^{2}$. For each initial condition, the ODEs are time integrated using the RK4 scheme with $\Delta t=0.01$ until $t=100$. To strictly enforce the manifold constraint, we normalize the state at every integration step so that its Euclidean norm remains unity. In addition, to account for the periodic nature of the dynamics, we truncate each path after one period. In total, there are $1578088$ sample points that are available as training data. For validation, 12 trajectories of $301$ steps are generated with the initial conditions uniformly distributed on $S^{2}$. For testing, another $32000$ trajectories of $1000$ steps are generated with the initial conditions randomly sampled on $S^{2}$. Subsequently, we map these datasets to $n$-dimensional torus for training, validation, and testing.
This example uses the KRR models based on DM and RBF using the skip-connection form. The convergence study considers the number of training samples $N=\{1024,2048,4096,8192\}$. These training samples are randomly selected from the training set in a nested manner, such that the smaller dataset of samples is a subset of the larger dataset. For each training dataset and each model, we conduct the validation study and determine the hyperparameters that produce the minimum average RMSE over the entire prediction horizon in the validation dataset and compute the mean over the 12 validation trajectories. The detailed validation results are provided in SI A.1.
2.4 Lorenz-63 Dynamics
$\dot{x}$ $=\sigma(y-x)$ (6a)
$\dot{y}$ $=x(\rho-z)-y$ (6b)
$\dot{z}$ $=xy-\beta z,$ (6c)
where $\sigma=10$, $\rho=28$, $\beta=8/3$, and ${\bf x}=[x,y,z]$ is used as the states. The forward invariant set of the system is a strange attractor, of a fractal dimension of approximately 2.06, and the maximal Lyapunov exponent is $\Lambda\approx 0.91$. To obtain trajectory data, we integrate the Lorenz-63 system using the RK4 scheme with a step size of 0.01s from a randomly sampled initial condition. To ensure the data is on attractor, we discard the first $4000$ steps, and retain the remaining steps as the dataset. For training and validation, a single long trajectory of $2\times 10^{5}$ steps is generated. For testing, we generate another long trajectory of $12.5\times 10^{5}$ steps, and divide it into 500 non-overlapping segments of length $2500$ steps, where each segment is treated as one test trajectory. This data generation strategy is consistent with the one reported in [mandal2025learning].
This example uses the KRR models in skip-connection form. For the convergence analysis, we randomly sample 500 time series of length $N+2N_{v}$ from the training trajectory to obtain 500 distinct data subsets. In each subset, the first $N$ points are used for training and three overlapping trajectories of length $N_{v}$ are sampled from the remaining $2N_{v}$ points for validation. In validation, the VPT criterion of $\gamma=0.3$ is used. For all $N$’s, $N_{v}=1500$ is used; further increment does not significantly improve the performance. More details of the validation procedure are provided in SI A.2. Eventually, the above procedure yields 500 independently trained and validated models.
In the top panel of Fig. 2, typical model predictions are illustrated for $N=4096$, where DM clearly shows a longer VPT than RBF. Subsequently, for a more systematic analysis, each of the 500 models is used to compute the VPT’s of 500 test trajectories, of which the mean is recorded; this process results in a total of $500$ mean VPT values. Table 1 reports the means and standard deviations of the $500$ mean VPT’s for each $N$. As $N$ increases, both KRR models show improved mean VPT, and reduced variance. Furthermore, across all data sizes, DM consistently outperforms RBF in the mean VPT with lower variance, except $N=512$. In addition, we highlight that, in the small data size, e.g., $N=1024$, the mean VPT of DM reaches $13.18$ using only $N+2N_{v}=4024$ data points in total. Compared to the state-of-the-art random feature model for chaotic dynamics [mandal2025learning], which uses 50000 data points, DM achieves 1.16 longer VPT units using over 12 times fewer samples. A more detailed comparison of VPT distributions is provided in SI A.2.
Model $N=512$ $N=1024$ $N=2048$ $N=4096$
RBF $8.87\pm 1.71$ $11.20\pm 1.42$ $12.72\pm 0.92$ $13.47\pm 0.61$
2.5 Kuramoto-Sivashinsky Equation, Chaotic Dynamics
$u_{t}+uu_{s}+u_{ss}+\nu u_{ssss}=0,\quad s\in[0,L],\ t\geq 0.$
In this problem, we consider DM and RBF models using the direct form. For the convergence study, we follow a similar procedure in the Lorenz-63 case, but this time using 50 time series of length $N+2N_{v}$ from the training trajectory to obtain 500 distinct data subsets. We consider $N=\{2048,4096,8192,16384\}$ and $N_{v}=2000$, and use VPT criterion of $\gamma=0.5$ for validation. More details of validation are provided in SI A.3. Lastly, each of the 50 models is tested against 500 test trajectories, of which the mean VPT is recorded; this yields $50$ mean VPT’s.
Model $N=2048$ $N=4096$ $N=8192$ $N=16384$
RBF $0.79\pm 0.13$ $1.49\pm 0.21$ $2.72\pm 0.26$ $4.31\pm 0.33$
2.6 Kuramoto-Sivashinsky Equation, Travelling Dynamics
Here, six models are considered:
  1. 1.
  2. 2.
    RBF: The same as DM, except that the RBF kernel is used.
  3. 3.
    GMKRR [huang2025learning]: The model is still based on KRR but uses a geometrically constrained operator-valued kernel. For consistency with the other baseline models, a special data preprocessing step based on prior knowledge (see SI A.4) is used for this model, so that the fast and slow timescale dynamics are learned separately, and only the first 100 steps are needed for training.
  4. 4.
  5. 5.
  6. 6.
See SI A.4 for more details in data preprocessing, model training, and validation.
Figure 3 compares the model predictions. The baselines NODE and LDNet, that do not account for the exact intrinsic dimension of 2, perform the worst. In particular, LDNet incorrectly predicts the slow timescale, while NODE fails to predict a constant slow timescale; this is not surprising as the invariant set is not leveraged during prediction. The CANDyMan performs slightly better than NODE and LDNet, especially in capturing the correct slow timescale, but error accumulates over long-time prediction, likely due to the inaccuracies in the NN approximation. As GMKRR leverages the prior knowledge like the previous three baselines, its error is one order of magnitude lower than the three, identifying the correct phase. The enhancement has been attributed to the explicit incorporation of geometrical constraints, including the tangent space of the manifold, in the formulation. However, the performance of GMKRR is severely bottlenecked by the accuracy of the estimated tangent space. Lastly, both the KRR models, DM and RBF, greatly outperform all the previous baselines, where the errors are reduced by 5-6 orders of magnitude. Furthermore, DM outperforms RBF by one order of magnitude, which is attributed to the capability of DM to implicitly account for the geometrical constraints.
2.7 Pitch-Plunge Flat Plate
To generate the datasets, we perform direct numerical simulations of the incompressible Navier–Stokes equations using an immersed-boundary projection method [taira2007immersed, colonius2008fast]. The simulations are performed on a $599\times 299$ Cartesian grid, resulting in a state dimension of $n=179101$; details of simulation are provided in SI A.5. The training data includes $12$ trajectories, using $\phi_{h}$’s from $0^{\circ}$ to $330^{\circ}$ with a step size of $\Delta\phi_{h}=30^{\circ}$. The validation data includes $3$ trajectories using $\phi_{h}=\{45^{\circ},225^{\circ},315^{\circ}\}$. The test trajectory uses $\phi_{h}=135^{\circ}$. Each trajectory is generated using a step size of 0.01s and then downsampled at every 10th data points; then the first 20s of trajectory is discarded to remove the transient response. Eventually, $200$ steps are generated for each of the training and validation trajectories, and $801$ steps are generated for the test trajectory. The long test trajectory, consisting of approximately $20$ periods of flow oscillation, is designed to assess the long-term prediction capability of the models.
This example uses the DM and RBF models using the direct form, with comparison to the ResDMD method. Due to the high-dimensionality of the problem, data pre-processing based on PCA is applied to reduce the dimension for the sake of computational efficiency. We choose $r=81$ PCA modes for the DM and RBF models, which retains 99.199.1% energy of the data. More modes, $r=162$, are retained for ResDMD, to mitigate its inherent spectral pollution issue. The model training, validation, and prediction are performed in the reduced-dimension space, and full-dimensional flow field is reconstructed from the model prediction. More details in data pre-processing, training, and validation are provided in SI A.5.
The typical snapshots of the model predictions and NRMSE’s are shown in the middle two rows of Fig. 4 for the test trajectory. The NRMSE at step $i$ shown in the third row is defined as $\text{NRMSE}={\|\hat{{\bf x}}_{i}-{\bf x}_{i}\|_{2}}/{\|{\bf x}_{i}\|_{2}}$, where the norm is over the space. Clearly, the RBF fails to capture the system dynamics and the vortex shedding phenomenon is completely missing. In contrast, DM and ResDMD provide relatively accurate predictions, where DM has the lower error. The last panel of Fig. 4 shows the temporal evolution of the error over the entire test prediction horizon. Again RBF deviates substantially from the truth even in the beginning of the horizon. There is a noticeable growth of error in the ResDMD model; this is due to the inaccurate eigenvalues identified by the model. In contrast, DM not only achieves the lowest overall error among the three models, but also consistently maintains the error level over the horizon, indicating that it is respecting the geometrical constraint over a long-term prediction of $20$ cycles. For more detailed comparisons, see SI A.5.
3 Conclusion and Discussion
4 Methods

### v2

@@ Abstract
In this work, we propose a simple kernel ridge regression (KRR) framework with a dynamic-aware validation strategy for long-term prediction of complex dynamical systems. By employing a data-driven kernel derived from diffusion maps, the proposed Diffusion Maps Kernel Ridge Regression (DM-KRR) method implicitly adapts to the intrinsic geometry of the system’s invariant set, without requiring explicit manifold reconstruction or attractor modeling, procedures that often limit predictive performance. Across a broad range of systems, including smooth manifolds, chaotic attractors, and high-dimensional spatiotemporal flows, DM-KRR consistently outperforms state-of-the-art random feature, neural-network and operator-learning methods in both accuracy and data efficiency. These findings underscore that long-term predictive skill depends not only on model expressiveness, but critically on respecting the geometric constraints encoded in the data through dynamically consistent model selection. Together, simplicity, geometry awareness, and strong empirical performance point to a promising path for reliable and efficient learning of complex dynamical systems.
@@ 1 Introduction
Broadly speaking, there are two common approaches in literature. The fully data-driven approach ignores the first-principle model and learns the entire solution operator or the vector field of the ODEs directly from observed data (see, e.g., [wang2005gaussian, williams2015data, brunton2016discovering, vlachas2018data, pathak2018model, hamzi2021learning, huang2025learning]). In contrast, the partially data-driven approach retains the first-principles model and learns only the model-error components that correct the underlying dynamics (see, e.g.,[kaheman2019learning, lin2021data, harlim2021machine, levine2022framework, venegas2025a]). In this study, we focus on the fully data-driven setting, although the proposed framework could, in principle, also be adapted for learning model-error components in partially data-driven formulations.
@@
Many methods have been proposed to identify low-dimensional representations of dynamical systems, ranging from linear techniques such as PCA (Principal Component Analysis) [jolliffe2011principal] and DMD (Dynamic Mode Decomposition) [williams2015data] and its variants, to nonlinear approaches based on autoencoders [Linot2020, Maulik2021, vlachas2018data, Regazzoni2024]. A limitation of PCA and DMD is that they sometimes fail to produce a finite, low-dimensional representation capable of capturing the full dynamics, even when the underlying invariant set itself is finite-dimensional. Moreover, algorithms in the DMD family approximate the Koopman operator and therefore require additional treatment to avoid spectral pollution. A promising algorithm to avoid spectral pollution in the Koopman spectral approximation is ResDMD [colbrook2024rigorous], which we include as a baseline in one of our test cases. As for autoencoders, it is found that the latent space only provides a global parametrization of the invariant set, whose dimension is typically larger than the intrinsic dimension of the invariant set, and the dynamics are not guaranteed to learn the exponential maps (for discrete-time model) or vector fields (for continuous-time model) while satisfying the invariant set constraint [huang2025learning]. As a result, the predicted trajectories may deviate from the invariant set, limiting the long-term prediction accuracy, as we shall see in several numerical examples with neural network (NN) baselines. Meanwhile, there is indeed an autoencoder-based approach, CANDyMan, that can uncover the invariant set when it is a smooth manifold [Floryan2022]. In such a case, the estimated dynamics are guaranteed to evolve on the manifold. However, this approach cannot be extended to the case where the invariant set is fractal. In addition, common to NN-based approaches, the computational effort to train the models is not negligible and the hyperparameter tuning process can be difficult.
@@
A promising approach with solid theoretical foundations for mitigating the curse of dimensionality is to exploit the “kernel trick,” which circumvents the explicit evaluation of high-dimensional feature spaces. For dynamical systems evolving on smooth manifolds (i.e., those whose forward invariant sets form smooth submanifolds of the phase space), the Geometric Constrained Multivariate Kernel Ridge Regression (GMKRR) method was recently introduced to approximate the governing vector field of the ODEs [huang2025learning]. GMKRR is paired with a Normal-Correction integration scheme that ensures time integration remains consistent with the manifold constraint. Benchmark studies have shown that GMKRR outperforms several state-of-the-art NN-based methods [Floryan2022, Regazzoni2024], achieving superior long-term predictive accuracy while reducing training costs by 3–4 orders of magnitude. Despite these promising results, GMKRR remains limited by the accuracy of tangent-bundle approximation and cannot be directly applied to systems whose invariant sets exhibit fractal structure.
@@
$x_{j}(t_{i+1})=\varphi_{j,\Delta t}({\bf x}(t_{i}))\approx f_{j}({\bf x}(t_{i}))=k_{\epsilon,N}({\bf x}_{i},X)\boldsymbol{\alpha}_{j}$ (1)
@@
$x_{j}(t_{i+1})-x_{j}(t_{i})=\varphi_{j,\Delta t}({\bf x}(t_{i}))-x_{j}(t_{i})\approx f_{j}({\bf x}(t_{i}))=k_{\epsilon,N}({\bf x}_{i},X)\boldsymbol{\alpha}_{j},$ (2)
@@
for each component $j=1,\ldots,n$. Here, $k_{\epsilon,N}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ denotes the data-dependent DM kernel with a lengthscale parameter $\epsilon$, $X=\{{\bf x}_{i}={\bf x}(t_{i})=(x_{1}(t_{i}),\ldots,x_{n}(t_{i}))\}_{i=1,\ldots,N}$ denotes the training dataset with $t_{i}=i\Delta t$, and $\boldsymbol{\alpha}_{j}$ denotes the $N$-dimensional regression coefficient for the $j$th component. While direct estimator can be used in general, we notice substantial improvement with skip-connection estimator when the underlying ODE systems are not numerically stiff. In this study, we consider a Kernel Ridge Regression solution with DM kernel, coined as Diffusion Maps Kernel Ridge Regression (DM-KRR), where the regression coefficient, $\boldsymbol{\alpha}_{j}=\boldsymbol{\alpha}_{j}(\epsilon,{\lambda_{\mathrm{reg}}})$, depends on the shape parameter, $\epsilon$, and ridge regression parameter, $\lambda_{\mathrm{reg}}$}. We will show that careful choice of validation metric to determine the model hyperparameters $(\epsilon,{\lambda_{\mathrm{reg}}})$ is essential for achieving optimal predictive performance. While the conclusion here resonates with the conclusion from the Kernel Flow (KF) approach [hamzi2021learning], the present approach differs significantly from KF. Particularly, KF parameterizes the flow map using a family of kernels and selects among them via validation over the same loss-square error, we employ a fixed, data-driven kernel derived the from diffusion maps and validate by comparing the prediction error of iterating the approximate model over a time window.
@@
Beyond its simplicity and strong empirical performance, our use of kernel methods is also theoretically motivated. If the underlying dynamical system is ergodic with invariant measure $\mu$ supported on a compact set $A\subset\mathcal{M}\subset\mathbb{R}^{n}$, then the Reproducing Kernel Hilbert Space (RKHS) induced by any continuous kernel $k$ on a compact set $A$ is universal for approximating $L^{2}(\mu)$ functions, even when $A$ lacks smoothness. Thus, in principle, one may employ any such kernel, e.g., Gaussian or Matérn. Additionally, when the forward invariant set $\mathcal{M}\supset A$ is a smooth $d$-dimensional sub-Riemannian manifold and the invariant measure is uniformly distributed with a compact support that lies in the invariant set, the diffusion maps (DM) kernel is a natural choice as it approximates the heat kernel in $L^{\infty}$-sense over the dataset as shown in [dunson2021spectral]. In effect, the DM kernel encodes the intrinsic geometry of the invariant set, allowing the learned dynamics to adapt seamlessly to its true intrinsic dimensionality.
The remainder of this paper is organized as follows. In Section 2, we give a quick overview of the kernel ridge regression for our setup and introduce a dynamic-aware validation strategy to select the model hyperparameters. In Section 3, we introduce the diffusion maps kernel and discuss some theoretical results that motivate this choice of kernel. In Section 4, we demonstrate the numerical performance of DM-KRR over a range of dynamical systems whose forward invariant sets may be smooth or non-smooth and are embedded in ambient spaces ranging from dimension $3$ to $10^{5}$. In Section 5, we close the paper with a short summary and discussion. We supplement the paper with four appendices. Appendix A gives a heuristic estimate to assist the validation procedure. The remaining three appendices (B-D) document additional details that correspond to the numerical examples.
2 Kernel Ridge Regression with Dynamic-Aware Validation Strategy
In this section, we give a short overview of kernel ridge regression (KRR) and a dynamic-aware validation strategy to select the model hyperparameters.